# Learning to Play Piano in the Real World

### v2

@@ III Learning to Play Piano in the Real World
Reward Design. We will briefly discuss each component of our reward function $r_{\text{total}}=r_{\text{energy}}+r_{\text{hand\_position}}+r_{\text{keypress}}+r_{\text{sliding}}$ in this section. However, a more detailed discussion is included in Appendix VI-A. The $r_{\text{energy}}$ reward penalizes fast movements that need a lot of force. Further, it avoids oscillations that don’t lead to higher rewards. The $r_{\text{sliding}}$ reward is designed to penalize fast sideways movement of the hand while keys are pressed to avoid damaging the robot. The reward $r_{\text{hand\_position}}$ is designed to guide the wrist towards intended keys, while $r_{\text{keypress}}$ encourages the agent to press the correct keys.
Evaluation Metrics. To quantify the success of the agent, we use the same metric as in the RoboPianist [17]. At each timestep, the current states of all piano keys are compared to the target states, addressing two questions:
  • ”How reliable is the model at not pressing the keys it should not press?” (Precision) and
  • ”How reliable is the model at pressing the keys it should press?” (Recall).
Both, the precision and recall score, range, from 0 to $1$. A score of $1$ indicates that the agent performed perfectly in this metric at all timesteps. Both metrics are combined using the harmonic mean to calculate the F1 score.
$F_{1}=\frac{2}{\text{recall}^{-1}+\text{precision}^{-1}}\,.$
We chose the F1 score as our primary score to quantify the success of the agent. Our driving argument is that we support the empirical statement, which was already made in the RoboPianist [17], that the F1 score represents the qualitative performance of the agent. Further, using the same metric across contributions makes the results more comparable and increases the transparency.
@@
Model Training. Our model consists of three hidden layers with 256 nodes each. The training uses DroQ [2], a soft actor-critic algorithm and runs for $1e6$ episodes. To reduce the Sim2Real gap, we utilized domain randomization (DR) of the height of the piano, the starting position of the hand, the dampening of the joints, the stiffness of the joints, the threshold of a key to be considered pressed, the stiffness of the springs holding the keys and the friction between the keys and the fingertips. Each simulation timestep is $0.005\text{\,}\mathrm{s}$. The current state is observed every 10th iteration, which leads to a control frequency of the policy of $20\text{\,}\mathrm{Hz}$.
@@
IV Experimental Results
IV-A How does our model perform? (Fig. 4(a))
Setup. To evaluate the performance of our approach, we trained on four different songs. Each of those songs was trained twice on independent seeds, which led to eight models in total. Since every run in the real world differs, we executed each model three times. The figure shows the average and standard deviation of the precision, recall and the resulting F1 score for each song after using hybrid execution as introduced in Section 3. Other execution modes are evaluated in Section IV-B. As a reference for the Sim2Real gap is also the performance in simulation provided.
@@ IV Experimental Results
Conclusion. In this experiment, we demonstrated that it is possible to train a piano-playing model using common RL techniques that can be deployed using hybrid execution. Our models are able to play all four songs successfully in simulation and in the real world. This demonstrates effectively that playing piano in the real world with a learned model is within the reach of our current capabilities.
IV-B Which execution mode has the most potential? (Fig. 4(b))
@@
Setup. In this experiment we perform an ablation study over the execution modes introduced in Section III: joint mirroring, hybrid execution, and real world execution. No DR was used for the joint mirroring and hybrid execution. For the real world execution we decided to use a moderate intensity of DR with $C_{dr}=0.5$. The effect of DR is further discussed in Section IV-C. Analog to the prior experiment, we trained each model configuration twice and executed each model 3 times per execution mode. The figure shows the average and standard deviation of 6 runs for each mode.
@@
Results. Results show that the agent’s performance differs across the execution modes depending on the song. While joint mirroring performed among the best for Twinkle Twinkle Little Star and the Chord Progression, it performed worse than other modes for both scales. The real world execution, on the other side, performed well for the C-Major Scale but underperformed for all other songs. The hybrid execution was the best-performing mode for three out of four songs. The only exception is Twinkle Twinkle Little Star, where the hybrid execution mode was slightly behind joint mirroring. To understand the superiority of hybrid execution, we need to discuss the advantages and disadvantages of each execution mode separately. When using joint mirroring, we don’t observe the real world, which makes it impossible to correct wrong keypresses. Further, the simulation might diverge from the real world which would reduce the performance. The advantage of this mode is that it does not need DR, which reduces the resource requirements and the number of hyperparameters. The lack of adaptability is especially a problem for both scales since the model learned to slide sideways across the keys during the training. Unfortunately, this strategy is hard to execute in the real world since the finger slips differently every time. The real world execution has the advantage of observing the whole state of the real world robot. This ideally allows the robot to react in the best way possible since no information is hidden. Unfortunately, the additional information also requires a more robust model, which can be reached by utilizing DR as shown in Section IV-C. However, DR makes the training environment significantly more demanding. In our experiments, the real world execution mode was not able to outperform the other execution modes. The hybrid execution combines both modes. Most of the observations are still generated in simulation, which makes it easier for the agent to transfer the learned knowledge into the real world. However, by providing observations of the actually pressed keys, the robot is able to correct its behavior. This way, the agent behaves similarly to joint mirroring when the pressed keys match the currently pressed keys in the simulation. However, as soon as the simulation differs from the real world, the agent is able to correct itself. This is possible while not using any DR at all, which makes the process even more favorable.
@@
Conclusion. We showed that hybrid execution performed the best in our setup. It allows the agent to correct itself while being able to use the joints in the same way as they behaved during training. This makes hybrid execution modes especially interesting for future work with the goal of achieving human-like proficiency in complicated settings, where it is hard to deploy a model entirely into the real world.
IV-C How does the intensity of DR affect the real world performance? (Fig. 5(a), Fig. 5(b))
@@
Conclusion. The best performance can be achieved with moderate DR. Too little DR leads to worse performance since the model is not able to recover successfully after the environment behaved in a different way compared to the training - the model is not robust enough. Too much DR on the other side worsens the performance since the model is not able to learn successfully during training. As a consequence of these results, we recommend researchers to carefully tune the intensity of DR, since it significantly impacts the real world performance.
V Limitations and Future Work
In this work, we share our findings on how to teach a robot to play simple songs on a real piano. We demonstrated a proof of concept that utilized Sim2Real and a hybrid execution mode to play piano in the real world. However, our system possesses several limitations that could be investigated in future research:
Song Generalization. Similarly to the RoboPianist [17], we observed that the performance decreases heavily when we attempted to test it on a different song than the one it was trained on. This means that we need to retrain the policy in simulation for each new song that we want to play. This is undesirable, as we would instead prefer to train a single policy that at execution time is conditioned on a set of tablatures. However, preliminary results show that naively training on multiple songs does not seem to solve this issue. Future research should investigate how to achieve this contextual policy in a way that is robust and efficient.
Tactile Sensors. Another limitation of our work is the lack of additional sensing modalities, including touch sensing. In prior works [16], it has been shown that touch can enable the robot to perceive additional information about the environment and improve performance. We considered integrating existing touch sensing solutions, such as DIGIT [4]. However, these sensors are larger than a human finger, and resulted in pressing two keys at once. We believe that integrating future generations of tactile sensors that are smaller and omnidirectional would be valuable and an interesting research topic.
Dexterous Hands. Existing multi-finger robotic hands are still not as dexterous as a human hand. This is particularly evident in the limitations of finger abduction and flexion. Moreover, the use of the wrist seems to be rather important in humans. These hardware limitations resulted in the excessive movement of the whole hand at the level of the arm. It would be interesting to study in the future how to reduce these movements to obtain more natural movement and exploit future generations of more dexterous hands.
Sliding Fingers. The current 3D-printed fingertip used in our experiments is made from PLA which is very stiff. This reduces the contact area of the fingers, which reduces the friction, which leads to sliding on the keys while playing a song. This makes playing the piano more demanding since the agent needs to balance the forces more precisely. Therefore, using fingertips made of softer materials, like TPU, silicone, or rubber, might increase the performance.
Bimanual Playing. More complex piano songs require the use of two hands. This also introduces the challenge of coordinating twice the number of joints and enables more complex fingering techniques. However, currently our hardware setup is limited to a single hand. In the future, we hope to extend to the bimanual case and investigate how to learn more complex piano movements, thus being able to increase the overall difficulty of the played songs.
VI Conclusion
In this work, we showcase the first proof-of-concept of how to learn to play piano with a real world robotic hand. Specifically, we employ a Sim2Real approach where we first train a policy in simulation, and then deploy it in the real world. We also introduce Hybrid Execution, a novel approach to infer with the model in the real world. Hybrid execution allows the agent to observe and act in the real world with a reduced Sim2Real gap compared to a traditional Sim2Real transfer. Hybrid execution has the most potential in complicated setups like piano-playing, where it is hard to obtain a model robust enough to be deployed traditionally using Real World Execution. Our experimental results demonstrate that our model can play several pieces successfully in the real world. Nonetheless, there is still a long way before we can achieve in robots the same level of proficiency and elegance as human pianists. Beyond the artistic raison d’être, we believe that playing piano is a relevant and challenging benchmark for dexterous contact-reach manipulation, and advances in this task can be more widely relevant for the whole manipulation community.
Acknowledgments
We thank Ken Nakahara for the support and tips regarding operating the robotics. This work was partly supported by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence “Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden, and by Bundesministerium für Bildung und Forschung (BMBF) and German Academic Exchange Service (DAAD) in project 57616814 (SECAI, School of Embedded and Composite AI).
Appendix
VI-A The Reward Formulation
The total reward consists of four parts
$r_{\text{total}}=r_{\text{energy}}+r_{\text{hand\_position}}+r_{\text{keypress}}+r_{\text{sliding}}\,$,
each rewarding a certain behavior of the agent.
The $r_{\text{energy}}$ reward penalizes fast movements that need a lot of force. Further, it avoids oscillations of fingers that don’t lead to higher rewards. The energy reward also leads to a policy that behaves more fluently, which makes it more human-like.
$r_{\text{energy}}=-|\tau_{\text{joints}}|^{T}|v_{\text{joints}}|*C_{\text{energy}}$
The implementation is the same as used in the RoboPianist environment [17], where $\tau_{\text{joints}}$ contains the torques of each active joint and $v_{\text{joints}}$ contains the velocities of each joint. $C_{\text{energy}}$ is the energy coefficient and is a hyperparameter of the environment. Unless otherwise specified $C_{\text{energy}}=0.12$.
The reward $r_{\text{hand\_position}}$ is designed to guide the hand towards targeted keys. This is achieved by reducing the average distance between the palm of the hand and the targeted keys.
[…]
The distances are bounded between 0 and $1$ using the tolerance function of the dm_control library [13], notated as $g$. Further is $P_{\text{targets}}$, the set of the cartesian positions of all targeted keys, and $p_{\text{palm}}$ represents the cartesian position of the palm.
The $r_{\text{keypress}}$ reward is the most important part of $r_{\text{total}}$ because it encourages the agent to press the correct keys. The reward is designed based on two design requirements:
  1. 1.
    Pressing no keys should be worse than pressing the wrong keys.
  2. 2.
    Pressing the correct keys should be better than pressing the wrong keys.
Those requirements lead the exploration of the model towards pressing the correct keys without being ”afraid” of pressing the wrong keys. This relationship is implemented by using multiple cases depending on the currently pressed keys:
For $k_{\text{target}}>0$ we divide the keypress reward into three cases
$r_{\text{keypress}}=\left\{\begin{aligned} &0&&,k_{\text{wrong}}=0\land k_{\text{correct}}=0&&\text{(1)}\\ &0.5+0.5*\bar{\mu}_{\text{target}}&&,k_{\text{wrong}}>0&&\text{(2)}\\ &1+0.5*\bar{\mu}_{\text{target}}&&,k_{\text{wrong}}=0\land k_{\text{correct}}>0&&\text{(3)}\end{aligned}\right$.
For $k_{\text{target}}=0$ is the reward calculated as follows:
$r_{\text{keypress}}=2*(1-\max_{\mu^{\prime}\in\mu_{\text{wrong}}}\mu^{\prime})$ (4)
The cases $(1)$, $(2)$ and $(3)$ assume that at least one key should be pressed at the current timestep, while the case $(4)$ defines the reward when no key is meant to be pressed. The cases are separated based on the currently pressed keys of the piano. For this purpose, the variable $k$ refers to a number of keys. More precisely:
  • $k_{\text{target}}$: How many keys are meant to be pressed?
  • $k_{\text{wrong}}$: How many keys are pressed that are not meant to be pressed?
  • $k_{\text{correct}}$: How many keys are pressed that are meant to be pressed?
Further, $\mu\in[0,1]$ refers to the normalized state of the key with $\mu=1$ representing a completely pressed key.
  • $\bar{\mu}_{\text{target}}$: The average state of all targeted keys.
  • $\mu_{\text{wrong}}$: The set of the states of all the pressed keys that are not meant to be pressed. This also implies that $|\mu_{\text{wrong}}|=k_{\text{wrong}}$.
Relating the cases to the design requirements reveals that case $(2)$ should yield a greater reward than case $(1)$. This condition is satisfied, as case $(2)$ always yields a positive value, whereas case $(1)$ is always zero. The second design requirement specifies that case $(3)$ should yield more reward than case $(2)$. This can be demonstrated by calculating the difference between the two cases: $(3)-(2)=1+0.5*\bar{\mu}_{\text{target}}-(0.5+0.5*\bar{\mu}_{\text{target}})=0.5>0$. Since this difference is always positive, case $(3)$ indeed provides a greater reward than case $(2)$, meeting the design requirement.
The $r_{\text{sliding}}$ reward is designed to penalize fast sideways movement of the hand while keys are pressed. This avoids the local optimum where one finger is pressed statically while the hand slider ”selects” the key. The penalty is doubled if two adjacent keys are pressed, which is unavoidable while sliding sideways.
$r_{\text{sliding}}=-\lambda*v_{\text{hand}}^{2}*3$
The $\lambda\in\{0,1,2\}$ represents how many adjacent keys are pressed simultaneously. $\lambda=0$ is the case if no keys are pressed, $\lambda=1$ is the case when no adjacent keys are pressed and $\lambda=2$ is the case when adjacent keys are pressed. Note that the current speed of the hand $v_{\text{hand}}$ is squared. This allows for slow hand movements, while fast movements are more strongly penalized.
The values of $\lambda\in\{0,1,2\}$ indicate the number of adjacent keys pressed simultaneously. Specifically, $\lambda=0$ represents no pressed keys, $\lambda=1$ corresponds to non-adjacent keys being pressed, and $\lambda=2$ denotes that adjacent keys are pressed. Additionally, the current hand speed $v_{\text{hand}}$ hand is squared, allowing slower hand movements to receive less penalty, while faster movements receive a stronger penalty. The whole term is multiplied by three to increase the effect of this penalty.

### v3

@@ Appendix A Our Hybrid Execution Setup
Reward Design. The reward function used in the following experiments $r_{\text{total}}=r_{\text{energy}}+r_{\text{hand\_position}}+r_{\text{keypress}}+r_{\text{sliding}}$ does not utilize optimal transport to explicitly encourage fingering. Instead, the reward $r_{\text{hand\_position}}$ is designed to guide the wrist towards intended keys, while $r_{\text{keypress}}$ encourages the agent to press the correct keys. The $r_{\text{sliding}}$ reward is designed to penalize fast sideways movement of the hand while keys are pressed to avoid damaging the robot. The $r_{\text{energy}}$ reward penalizes fast movements that need a lot of force. Further, it avoids oscillations that don’t lead to higher rewards.
@@
Model Training. During training we use domain randomization (DR) as introduced in Section III. However we don’t apply our Sim2Real2Sim procedure. Instead, we focus on three execution modes.
@@
Appendix B Preliminary Results
We provide videos showcasing real world runs for each of the experiments on our website: www.lasr.org/research/learning-to-play-piano. In this section, we experimentally answer the following questions:
B-A How does our model perform? (Fig. 9(a))
Setup. To evaluate the performance of our approach, we trained on four different songs. Each of those songs was trained twice on independent seeds, which led to eight models in total. Since every run in the real world differs, we executed each model three times. The figure shows the average and standard deviation of the precision, recall and the resulting F1 score for each song after using hybrid execution as introduced in Appendix 8. Other execution modes are evaluated in Appendix B-B. As a reference for the Sim2Real gap is also the performance in simulation provided.
@@ Appendix B Preliminary Results
Discussion. In this experiment, we demonstrated that it is possible to train a piano-playing model using common RL techniques that can be deployed using hybrid execution. Our models are able to play all four songs successfully in simulation and in the real world. This demonstrates effectively that playing piano in the real world with a learned model is within the reach of our current capabilities.
B-B Which execution mode has the most potential? (Fig. 9(b))
@@
Setup. In this experiment we perform an ablation study over the execution modes introduced in Appendix C: joint mirroring, hybrid execution, and real world execution. No DR was used for the joint mirroring and hybrid execution. For the real world execution we decided to use a moderate intensity of DR with $C_{dr}=0.5$. The effect of DR is further discussed in Appendix B-C. Analog to the prior experiment, we trained each model configuration twice and executed each model 3 times per execution mode. The figure shows the average and standard deviation of 6 runs for each mode.
@@
Results. Results show that the agent’s performance differs across the execution modes depending on the song. While joint mirroring performed among the best for Twinkle Twinkle Little Star and the Chord Progression, it performed worse than other modes for both scales. The real world execution, on the other side, performed well for the C-Major Scale but underperformed for all other songs. The hybrid execution was the best-performing mode for three out of four songs. The only exception is Twinkle Twinkle Little Star, where the hybrid execution mode was slightly behind joint mirroring. To understand the superiority of hybrid execution, we need to discuss the advantages and disadvantages of each execution mode separately. When using joint mirroring, we don’t observe the real world, which makes it impossible to correct wrong keypresses. Further, the simulation might diverge from the real world which would reduce the performance. The advantage of this mode is that it does not need DR, which reduces the resource requirements and the number of hyperparameters. The lack of adaptability is especially a problem for both scales since the model learned to slide sideways across the keys during the training. Unfortunately, this strategy is hard to execute in the real world since the finger slips differently every time. The real world execution has the advantage of observing the whole state of the real world robot. This ideally allows the robot to react in the best way possible since no information is hidden. Unfortunately, the additional information also requires a more robust model, which can be reached by utilizing DR as shown in Appendix B-C. However, DR makes the training environment significantly more demanding. In our experiments, the real world execution mode was not able to outperform the other execution modes. The hybrid execution combines both modes. Most of the observations are still generated in simulation, which makes it easier for the agent to transfer the learned knowledge into the real world. However, by providing observations of the actually pressed keys, the robot is able to correct its behavior. This way, the agent behaves similarly to joint mirroring when the pressed keys match the currently pressed keys in the simulation. However, as soon as the simulation differs from the real world, the agent is able to correct itself. This is possible while not using any DR at all, which makes the process even more favorable.
@@
Discussion. We showed that hybrid execution performed the best in our setup. It allows the agent to correct itself while being able to use the joints in the same way as they behaved during training. This makes hybrid execution modes especially interesting for future work with the goal of achieving human-like proficiency in complicated settings, where it is hard to deploy a model entirely into the real world.
B-C How does the intensity of DR affect the real world performance? (Fig. 10(a), Fig. 10(b))
@@
Discussion. The best performance can be achieved with moderate DR. Too little DR leads to worse performance since the model is not able to recover successfully after the environment behaved in a different way compared to the training - the model is not robust enough. Too much DR on the other side worsens the performance since the model is not able to learn successfully during training. As a consequence of these results, we recommend researchers to carefully tune the intensity of DR, since it significantly impacts the real world performance.
Appendix C Conclusion of Hybrid Execution
In the appendix, we present an alternative approach for learning to play piano in the real world. Specifically, we adopt a Sim2Real paradigm, in which a policy is first trained in simulation and subsequently deployed in the real world environment. We further introduce Hybrid Execution, a novel inference framework designed for real world deployment. Hybrid Execution enables the agent to both observe and act in the real world with a reduced Sim2Real gap. Hybrid Execution is particularly well-suited for complex tasks such as piano playing, where achieving sufficient robustness for direct real-world deployment remains challenging. However, an important consideration for applying Hybrid Execution more broadly is the potential divergence between simulation and reality, which may lead to degraded performance. Overall, our experimental results demonstrate that Hybrid Execution enables the robotic system to successfully perform several simple piano pieces in a real-world setting.